A pharmacist has an 18% bleach solution. How much of this solution and how much water must be mixed together to make 10 liters of a 12% solution?
Answer by lwsshak3(11628)

A pharmacist has an 18% bleach solution. How much of this solution and how much water must be mixed together to make 10 liters of a 12% solution?
***
let x=amt of water to mix with bleach
10-x=amt of 18% bleach to mix with water
18%(10-x)=12%*10
1.8-.18x=1.2
.18x=0.6
x=3 1/3
10-x=6 2/3
amt of water to mix with bleach =3 1/3 liters
amt of 18% bleach to mix with water=6 2/3 liters